OpenStudy (anonymous):

Simplify the expression. sin^2 x-1/cos(-x)

OpenStudy (anonymous):

\[\sin^2x=?\]

OpenStudy (anonymous):

(Sin x)^2 perhaps

OpenStudy (anonymous):

\[\cos(-x)=\cos x\]

OpenStudy (anonymous):

\[\sin ^{2}x+\cos ^{2}x=1\] \[\sin ^{2}x=1-\cos ^{2}x\]

OpenStudy (anonymous):

\[\frac{ \sin ^{2}x-1 }{\cos(-x) }=\frac{ 1-\cos ^{2}x-1 }{\cos x}=\frac{ - \cos ^{2}x }{ \cos x }=-\cos x\]

OpenStudy (anonymous):

use identities